# Automated UAV-based Wind Turbine Blade Inspection: Blade Stop Angle Estimation and Blade Detail Prioritized Exposure Adjustment

### v1

@@ I INTRODUCTION
In recent years, propelled by the relentless progress in renewable energy technologies, wind power has emerged as a significant electricity source. As of 2023, the global installed wind power capacity surged to an impressive 1021GW [1]. Wind farms strategically dot hills, mountains, and coastal regions, leveraging their bountiful wind resources to offer substantial energy potential for efficient wind turbine installations [2]. Despite these advancements, severe weather conditions such as intense winds, lightning, hail, and rain pose threats to wind turbine blades, potentially leading to damage [3]. Regular inspections of these blades are vital for promptly identifying and rectifying any issues, thereby averting disruptions to the turbine’s normal operation.
@@ I-A Design of UAV Inspection Platform
Long Endurance: Inspections often target large wind turbines up to 130m high and 50m blades, necessitating a safety distance of over 10m from the blades, and a flight speed below 2m/s to prevent motion blur in images capture. Such long distances and low speeds demand long endurance.
@@ I-B Robust Blade Stop Angle Estimation
During inspection, the wind turbine rotor must be stopped by the braking system [14]. At this point, the angle between the turbine tower and the first blade to its right is defined as the blade stop angle. Given that wind turbine blades can only rotate passively by wind, it is time-consuming and unrealistic to stop the blades at a specified angle, as this would require waiting for the unpredictable wind to position the blades at that angle, especially in windless weather. Hence, the turbine blades may be positioned at any stop angle, result in uncertain blade orientation. By estimating the blade stop angle during inspection, it becomes possible to accurately ascertain the blade’s position and orientation, thereby aiding in subsequent inspection planning and control [7].
@@ III-A UAV Inspection Platform
To meet the demands from automated inspection tasks and scenarios, we propose a platform depicted in Fig. 2(a). The platform is based on the DJI M300 drone, with 30-minute long endurance, 12m/s wind resistance, and obstacle sensing capabilities, ensuring stable and secure inspections. The Livox MID-360 LiDAR, offering a 70m range for 3D perception, is well-suited for the large-scale point cloud acquisition required by perception algorithms. Considering the forward perception demand and the LiDAR’s field of view (FOV), the placement is illustrated in Fig. 2(b). The DJI H20T gimbal camera enables 20MP high-quality image capture, camera orientation control and exposure parameter adjustments. The Intel NUC11TNKi5 serves as the onboard computer, with a 4 cores and 8 threads 4.20GHz CPU and 64G memory, providing ample computing power. In addition, a protective shell was designed for the onboard computer using 3D printing to provide protection against rain and sand.
@@ III-B Fermat Point Based Blade Stop Angle Estimation
Motivated by the unique spatial geometric attributes of wind turbines, we design a blade stop angle estimation approach. As depicted in Fig. 3, envision the utilization of a ring, exceeding the hub’s size and parallel to the blade rotation plane, to intersect the 3 blades along the hub’s orientation. This intersection yields 3 Blade-Ring Intersection (BRI) regions on the blade segments, with their respective centers as BRI points. A triangle is formed with these 3 BRI points as the BRI triangle. The Fermat point of this triangle, which defined by Def. 1, can be found according to [19].
@@ III-B1 Blade Rotation Plane Extraction
As shown in Fig. 4(b), the drone hovers in front of the wind turbine, orienting towards the hub, with the position denoted as $\bm{p}_{d}$. $\bm{p}_{d}$ satisfies $\bm{p}_{d}=\bm{p}_{h}^{0}-\bm{v}_{n}\cdot d_{c}$, where $d_{c}$ is the desired distance between the drone and the hub, and $\bm{p}_{h}^{0}$ is an estimate of hub position with errors, obtained from [9, 12, 15] as the drone ascends along the tower. The point cloud of the wind turbine is captured by LiDAR, denoted as $\bm{\mathcal{P}}_{a}$. Plane fitting is applied on $\bm{\mathcal{P}}_{a}$ using the Random Sample Consensus algorithm [20], resulting in the blade rotation plane $\bm{\Pi}_{\bm{\mathcal{P}}}$ and its unit normal vector $\bm{v}_{n}$.
@@ III-B2 Ring Region Searching and Clustering
III-B3 Fermat Point Optimization
Finally, the hub center is optimized utilizing the Prop. 1 and the blade stop angle is computed, as shown in Fig. 4(g). A triangle $\triangle B$ is constructed using $\bm{Q}_{b}$. The Fermat point of $\triangle B$ is determined following [19] and is applied as the optimal estimation for the hub center $\bm{p}_{h}$. The hub center $\bm{p}_{h}$ and $\bm{Q}_{b}$ is connected to determine the blade direction vector $\bm{V}_{b}=\{\bm{v}_{b1},\bm{v}_{b2},\bm{v}_{b3}\}$ and compute the blade stop angle $a_{b}$. The algorithm’s convergence criterion is met when $a_{b}^{i}$ obtained in the $i$-iteration and $a_{b}^{i-1}$ acquired in the $(i-1)$-iteration satisfy $|a_{b}^{i}-a_{b}^{i-1}|<\varepsilon_{a}$, leading to the optimal $\hat{a}_{b}$ and $\hat{\bm{p}}_{h}$. The $\varepsilon_{a}$ denotes the convergence threshold. The algorithm procedure is summarized in Algorithm. 1.
@@ III-C Blade Details Prioritized Exposure Adjustment
To ensure clear details of the blade region in the captured image, adjusting the exposure based on the brightness of the blade region is necessary. For this purpose, we introduce a blade detail prioritized exposure adjustment approach. Considering that the point cloud provides a spatial representation of the blade, the proposed approach initially utilizes the point cloud to extract the current inspection point on the blade. Subsequently, this inspection point is projected onto the image, serving as the reference point for exposure adjustment. Finally, the camera exposure parameter is adjusted according to the brightness within the reference region surrounding the reference point. Further details are outlined below.
@@ IV EXPERIMENT
We conducted experiments in operational wind farms. To date, the proposed platform and approaches have successfully completed 120 flights involving 10 distinct models of wind turbines across 5 operational wind farms. Throughout these experiments, temperatures varied from 5C to 33C, with maximum wind speeds reaching up to $12$m/s. Despite an occasional interruption caused by sudden light rain, the platform safely returned without any damage. These results highlight the performance of our platform in coping with the complex environment of the wind farm. In the following, we will present the results of the blade stop angle estimation and the exposure adjustment approaches, respectively.
@@ IV-A Fermat Point Based Blade Stop Angle Estimation
Mean Angle Error $(^{\circ})$ $\bm{\downarrow}$ 2.14 1.15
@@
Figure 7: Result of the blade exposure adjustment algorithm: (a)-(d) represent the underexposure scenario, where (a)&(b) show the original and adjusted images, and (c)&(d) are the histograms of (a)&(b), respectively. (e)-(h) represent the overexposure scenario, where (e)&(f) show the original and adjusted images, and (g)&(h) are the histograms of (e)&(f), respectively. (i)&(j) present the metric curves of the original and adjusted images in the scenarios of underexposure and overexposure respectively, with the x-axis represents the distance from the drone and the blade root.
Tab. I outlines the mean angle error and success rates of both methods. The proposed method exhibits a mean angle error of 1.15, surpassing the 2.14 error of the vision-based method. Additionally, the proposed method boasts a success rate of 98.3%, significantly outperforming the 69.2% success rate of the vision-based method. These findings underscore the superior robustness and accuracy of the proposed method.
@@ IV-B Blade Details Prioritized Exposure Adjustment
Comparing the original and adjusted images, we analyze their histograms within the blade region. The position and width of the histogram illustrate brightness and detail richness. Underexposure, as shown in Fig. 7(a), manifests as dimness and invisible details (red circle in Fig. 7(a)&(b)). The adjusted Fig. 7(b) enhances brightness, revealing clearer details. In Fig. 7(c) and Fig. 7(d), the histogram transitions from a low gray value of 22.099 to 140.280, expanding the gray range by 96.96%. Conversely, overexposure, depicted in Fig. 7(e), leads to excessive brightness and blurry details (red circle in Fig. 7(e)&(f)). The adjusted Fig. 7(f) rectifies this, enhancing intricacy of details. In Fig. 7(g) and Fig. 7(h), the histogram shifts from a high gray value of 250.733 to of 135.595, expanding the gray range by 51.29%.
We compute these metrics for original and adjusted images, plotting their curves along the blade’s position. Fig. 7(i) and Fig. 7(j) illustrate underexposed and overexposed scenarios. The original $\mu$ curve lies at the extremes, while the adjusted curve settles within an optimal range, indicating appropriately adjusted brightness. The adjusted $\sigma$ and $H$ curves surpass the original, signifying richer blade details. Tab. II showcases the mean of these metrics. In the underexposed scenario, the blade brightness increased from a low value of 25.48 to 133.27, The richness of blade details, quantified by standard deviation and entropy, increased by 218.85% and 56.19% respectively. Conversely, in the overexposed scene, the blade brightness was adjusted from a high value of 245.64 to 150.85. The richness of blade details, measured in terms of standard deviation and entropy, saw increments of 187.73% and 40.65% respectively. These underscore the effectiveness of the proposed method in both two scenarios, ensuring that the brightness falls within an optimal range and preserving intricate details within blade region effectively.
@@
Std Dev $\sigma$ $\bm{\uparrow}$ 2.48 7.91
Entropy $H$ $\bm{\uparrow}$ 2.10 3.28
@@
Std Dev $\sigma$ $\bm{\uparrow}$ 5.46 10.25
Entropy $H$ $\bm{\uparrow}$ 1.23 2.96
@@ V CONCLUSION AND FUTURE WORK
In this paper, toward automated UAV-based wind turbine blade inspection, we propose a UAV inspection platform, a Fermat point based blade stop angle estimation approach, and a blade details prioritized exposure adjustment approach. Extensive 120 flights conducted across 10 wind turbine models in 5 operational wind farms have demonstrated their superior performance. The platform demonstrated stability and safety during all flights. The blade stop angle estimation approach achieves an angle error of 1.15 with a 98.3% success rate. Additionally, the exposure adjustment approach maintains optimal brightness within the blade region, enhancing the intricacy of details by 56.19% and 40.65% in underexposed and overexposed scenarios, respectively.

### v2

@@ I INTRODUCTION
In recent years, propelled by the relentless progress in renewable energy technologies, wind power has emerged as a significant electricity source. As of 2023, the global installed wind power capacity surged to an impressive 1021GW [1]. Wind farms strategically dot hills, mountains, and coastal regions, leveraging their bountiful wind resources to offer substantial wind energy potential to support efficient wind turbine installations [2]. Despite these developments in wind power technology and capacity growth, severe weather conditions such as intense winds, lightning, hail, and rain pose threats to wind turbine blades, potentially leading to damage [3]. Regular inspections of these blades are vital for promptly identifying and rectifying any issues, thereby averting disruptions to the turbine’s normal operation.
@@ I-A Design of UAV Inspection Platform
Long Endurance: Inspections often target large wind turbines, such as those up to 130 $\mathrm{m}$ high with 50 $\mathrm{m}$ blades, necessitating a safety distance of over 10m from the blades, and a flight speed below 2 $\mathrm{m/s}$ to prevent motion blur in image capture. Such long distances and low speeds demand long endurance.
@@ I-B Robust Blade Stop Angle Estimation
During inspection, the wind turbine rotor must be stopped by the braking system [14]. At this point, the angle between the turbine tower and the first blade to its right is defined as the blade stop angle. Given that wind turbine blades can only rotate passively by wind, it is time-consuming and unrealistic to stop the blades at a specified angle, as this would require waiting for the unpredictable wind to position the blades at that angle, especially in windless weather. Hence, the turbine blades may be positioned at any stop angle, resulting in uncertain blade orientation. By estimating the blade stop angle during inspection, it becomes possible to accurately ascertain the blade’s position and orientation, thereby aiding in subsequent inspection tasks. This aligns with [7, 8], which highlights that this angle is essential for effective path planning and trajectory generation in wind turbine inspection.
@@ III METHODOLOGY
This section details the methodologies developed to address the three core challenges in Sec. I: UAV platform design, robust blade stop angle estimation, and real-time blade exposure adjustment.
@@ III-A UAV Inspection Platform
To meet the demands from automated inspection tasks and scenarios mentioned in Sec. I-A, we propose a platform depicted in Fig. 2(a). The platform is based on the DJI M300 drone, with 30-minute long endurance, 12m/s wind resistance, and obstacle sensing capabilities, ensuring stable and secure inspections. The Livox MID-360 LiDAR, offering a 70m range for 3D perception, is well-suited for the large-scale point cloud acquisition required by perception algorithms. Considering the forward perception demand and the LiDAR’s field of view (FOV), the placement is illustrated in Fig. 2(b). The DJI H20T gimbal camera enables 20MP high-quality image capture, camera orientation control and exposure parameter adjustments. The Intel NUC11TNKi5 serves as the onboard computer, with a 4 cores and 8 threads 4.20GHz CPU and 64G memory, providing ample computing power. In addition, a protective shell was designed for the onboard computer using 3D printing to provide protection against rain and sand.
@@ III-B Fermat Point Based Blade Stop Angle Estimation
Motivated by the unique spatial geometric attributes of wind turbines, we design a blade stop angle estimation approach. This method targets the challenge of environment-sensitive angle estimation mentioned in Sec. II-B, leveraging the inherent 120 blade spacing to achieve robust and accurate estimation. The convergence threshold $\varepsilon_{a}$ is set to ensure the accuracy level required for reliable inspection planning.
As depicted in Fig. 3, envision the utilization of a ring, exceeding the hub’s size and parallel to the blade rotation plane, to intersect the 3 blades along the hub’s orientation. This intersection yields 3 Blade-Ring Intersection (BRI) regions on the blade segments, with their respective centers as BRI points. A triangle is formed with these 3 BRI points as the BRI triangle. The Fermat point of this triangle, which defined by Def. 1, can be found according to [19].
@@ III-B1 Blade Rotation Plane Extraction
As shown in Fig. 4(b), the drone hovers in front of the wind turbine, orienting towards the hub, with the position denoted as $\bm{p}_{d}$. $\bm{p}_{d}$ satisfies $\bm{p}_{d}=\bm{p}_{h}^{0}-\bm{v}_{n}\cdot d_{c}$, where $d_{c}$ is the desired distance between the drone and the hub, and $\bm{p}_{h}^{0}$ is a prior estimate of hub position with considerable errors, obtained from [9] as the drone ascends along the tower. The point cloud of the wind turbine is captured by LiDAR, denoted as $\bm{\mathcal{P}}_{a}$. Plane fitting is applied on $\bm{\mathcal{P}}_{a}$ using the Random Sample Consensus algorithm [20], resulting in the blade rotation plane $\bm{\Pi}_{\bm{\mathcal{P}}}$ and its unit normal vector $\bm{v}_{n}$.
@@ III-C Blade Details Prioritized Exposure Adjustment
To ensure clear details of the blade region in the captured images, we introduce a blade detail prioritized exposure adjustment approach. This approach addresses the gap in real-time blade-focused exposure mentioned in Sec. II-C, adjusting the camera exposure parameter to keep blade surface illumination within $\left[\mu_{\text{min}},\mu_{\text{max}}\right]$, which is derived from the specific requirements of subsequent damage detection tasks.
Considering that the point cloud provides a spatial representation of the blade, the proposed approach initially utilizes the point cloud to extract the current inspection point on the blade. Subsequently, this inspection point is projected onto the image, serving as the reference point for exposure adjustment. Finally, the camera exposure parameter is adjusted according to the brightness within the reference region surrounding the reference point. Further details are outlined below.
@@ IV EXPERIMENT
We conducted experiments in operational wind farms. To date, the proposed platform and approaches have successfully completed 120 flights involving 10 distinct models of wind turbines across 5 operational wind farms. Throughout these experiments, temperatures varied from 5C to 33C, with maximum wind speeds reaching up to 121212m/s. Despite an occasional interruption caused by sudden light rain, the platform safely returned without any damage. These results highlight the performance of our platform in coping with the complex environment of the wind farm. In the following, we will present the results of the blade stop angle estimation and the exposure adjustment approaches, respectively.
@@ IV-A Fermat Point Based Blade Stop Angle Estimation
Mean Angle Error $(^{\circ})$${\downarrow}$ 2.14 1.15
@@
Figure 7: Results of the blade exposure adjustment algorithm, showing two scenarios: (i) underexposure and (ii) overexposure. For both scenarios, columns (a) and (c) display the original and adjusted images, respectively, with their corresponding histograms in columns (b) and (d). Column (e) presents metric curves comparing the original and adjusted images across the inspection process of a single blade surface, where the x-axis represents the distance from the drone to the blade root. (Abbreviation: Ori.: Original, Adj.: Adjusted, Dist.: Distance)
Tab. I compares the mean angle error and success rates of the two methods. The proposed method outperforms the vision-based method, with a lower mean angle error (1.15 vs. 2.14) and a significantly higher success rate (98.3% vs. 69.2%). These findings underscore the superior robustness and accuracy of the proposed method. Among these 120 trials, 2 instances of failure were observed. One arose from heavy fog, which impaired LiDAR’s perception of blade structure. The other stemmed from large errors in the prior hub position estimate $\bm{p}_{h}^{0}$, leaving the ring region searching unable to extract the BRI regions. Future work will focus on improving the accuracy of the prior hub position estimate.
@@ IV-B Blade Details Prioritized Exposure Adjustment
Comparing the original and adjusted images, we analyze their histograms within the blade region. The position and width of the histogram reflect brightness and detail richness, respectively. Underexposure, as shown in Fig. 7(i)-(a), manifests as dimness and invisible details (red circle in Fig. 7(i)-(a)). The adjusted image Fig. 7(i)-(c) enhances brightness to reveal clearer details. In Fig. 7(i)-(b) and (i)-(d), the corresponding histograms show a transition from a low gray value of 22.10 to 140.28, expanding the gray range by 97.0%. Conversely, overexposure, depicted in Fig. 7(ii)-(a), leads to excessive brightness and blurry details (red circle in Fig. 7(ii)-(a)). The adjusted image Fig. 7(ii)-(c) rectifies this to enhance intricacy of details. In Fig. 7(ii)-(b) and (ii)-(d), their corresponding histograms shift from a high gray value of 250.73 to of 135.60, expanding the gray range by 51.3%.
We compute these metrics for both the original and adjusted images, plotting their curves as the drone inspects from blade root to tip. Fig. 7(i)-(e) and (ii)-(e) illustrate underexposed and overexposed scenarios, respectively. The original $\mu$ curve lies at the extremes, while the adjusted curve settles within an optimal range, indicating appropriately adjusted brightness. The adjusted $\sigma$ and $H$ curves surpass the original, signifying richer blade details. Tab. II showcases the metric means. In underexposed scenarios, the brightness increased from 25.48 to 133.27, with the detail richness quantified by $\sigma$ and $H$ improving by 218.85% and 56.19% respectively. In overexposed scenarios, the brightness was adjusted from 245.64 to 150.85, with detail richness up by 187.73% and 40.65% respectively. These results confirm the effectiveness of the proposed method in both scenarios, maintaining appropriate brightness and preserving intricate blade details. Suboptimal performance occasionally occurred during rapid sunlight changes. The slight delay in exposure adjustment led to temporary overexposure or underexposure in a few consecutive frames, indicating room to improve response speed for fast-changing lighting.
@@
Std Dev $\sigma$${\uparrow}$ 2.48 7.91
Entropy $H$${\uparrow}$ 2.10 3.28
@@
Std Dev $\sigma$${\uparrow}$ 5.46 10.25
Entropy $H$${\uparrow}$ 1.23 2.96
@@ V CONCLUSION AND FUTURE WORK
In this paper, toward automated UAV-based wind turbine blade inspection, we propose a UAV inspection platform, a Fermat point based blade stop angle estimation approach, and a blade details prioritized exposure adjustment approach. Extensive 120 flights conducted across 10 wind turbine models in 5 operational wind farms have demonstrated their superior performance, with tangible contributions to addressing the core challenges of automated inspection. The platform’s stability ensures consistent data collection in complex wind farm environments, addressing the operational reliability challenges. The blade stop angle estimation approach, with a mean error of 1.15 and a success rate of 98.3%, enhances blade position accuracy to facilitate more accurate inspection trajectory planning, and reduce flight tracking errors along the blade surface. This is critical for ensuring full coverage of blade inspection. Additionally, the exposure adjustment approach maintains optimal blade brightness and enhances detail clarity, with improvements of 56.19% and 40.65% in underexposed and overexposed scenarios. This makes subtle defects distinguishable, directly improving the reliability of downstream damage detection.